OpenStudy (anonymous):

If BC is parallel to DE, then all of the following are equivalent except *drawing inc* AD/AB DE/BC CE/AC AE/AC

OpenStudy (anonymous):



OpenStudy (anonymous):

CE/AC

OpenStudy (anonymous):

to be equivalent to the others it should be AC / CE

OpenStudy (anonymous):

same picture different question AC = 4, AE = 7, AD = 10, what is the length of AB? answers: 1 3/7 2 1/2 5 5/7 17 1/2

OpenStudy (anonymous):

4/7 = AB/ 10 AB = 40/7 = 5 5/7

OpenStudy (anonymous):

theres 2 more questions with that same picture..can u please answer them?

OpenStudy (anonymous):

i wont answer them but i'll point you in the right direction - i'd like you to be able to do them yourself

OpenStudy (anonymous):

ok..BC parallel to DE. BC = 6, DE = 9 Which of the following could be the lengths of AB and BD respectively?

OpenStudy (anonymous):

the large and small triangles are similar so corresponding sides are in same proportion. so BC/DE = 6/9 = AB / AD and BD = AD- AB

OpenStudy (anonymous):

so whats the answer